"Florida Citrus Inc. (FCI) estimates its taxable income at $7,100,000. Use the Corporate Tax Schedule, Table 9.12, to calculate how much FCI will pay in federal income taxes."


TABLE 9·12 Corporate Tax Schedule for 2015 Taxable Income (X) $0-$50,000 50,001-75,000 75,001-100,000 100,001-335,000 335,001-10,000,000 10,000,001-15,000,000 15,000,001-18,333,333 18,333,334 and up Tax Rate 15% 25% 34% 34% + 5% 34% 35% 35% + 3% 35% Tax Computation Formula $0 + 0.15X 7,500 + 0.25(X - $50,000) 13,750 + 0.34(X - 75,000) 22,250 + 0.39(X - 100,000) 113,900 +0.34(X- 335,000) 3,400,000 + 0.35(X - 10,000,000) 5,150,000 + 0.38(X - 15,000,000) 6,416,666 + 0.35(X - 18,333,333)
Answer #1

Now X=$7,000,000

335,01<7,000,000<10,000000

therefore appropriate tax to be paid will be 113900+0.34(X-335000)=113900+0.34(7000000-335000)=$2,380,000